Consider the following statistics from a Labor Market and then answer the questions which follow;

• 8,570 workers have been laid-off from employment and are waiting to be recalled.

• 4,523 have lost morale and quit actively seeking employment.

• 3,498 people are below 16years.

• 53,862 people are currently in education and not searching for employment.

• 9,454 workers have been retired from employment.

• 57,367 workers are employed but 65% are still searching for better employment.

• 378 people have switched between jobs.

• 157 quit their jobs because of low salaries and are seeking better paying jobs.

• 12,321 new graduates are actively seeking formal employment.

• 1,563 are in jail.

a) How many people are discouraged workers?

b) How many people are unemployed?

c) How many people are in the Labor Force?

d) Calculate the unemployment rate?

e) How many people are not in the Labor Force?

f) How many people are part of the Adult Working Population?




1
Expert's answer
2021-06-21T12:05:11-0400

a)

Discouraged workers are those who lost their morale and quit actively seeking job employment. Number of discouraged labor is 4523.


b)

Unemployed population:- 

Unemployed labor force= 12321+8570+157

                    = 21048


C)

Labor Force:- It includes total number of people who are currently employed and those who are unemployed and seeking employment.  

Labor force= Employed + Unemployed force

          = 57367+21048

          = 78415


d)

"Unemployment\space rate=\frac{unemployed population}{total\space labor\space force}\times100"


here; 

Unemployed population= 21048

Total labor force= 78415

therefore;


"Unemployment\space rate=\frac{21048}{78415}\times100"


"=26.84\%"


e)

People who are not in the labor force includes:- People who are below 16 years, discouraged labor, retired employees, people who are currently in education and not searching for employment and people in jail 

Not in labor force= 3498+53862+9454+4523+1563

              = 72900


f)

Adult working population= Labor force+ Man labor force

                     = 78415+69402

                     = 148195